# FlashVault: Versatile In-NAND Self-Encryption with Zero Area Overhead

### v1

@@ III-B2 Performance of Block Cipher Algorithms
Fig. 5 shows the latency breakdown for six representative block ciphers (3DES [84], HIGHT [37], IDEA [59], Serpent [9], SM4 [33], Camellia [4]) executed in CTR mode, measured under both CPU and FlashVault. Although FlashVault embeds multiple cryptographic engines within each NAND flash chip, Fig. 5 reports the result based on a single-block-per-chip configuration to focus on the impact of where encryption occurs. The experimental setup for FlashVault is described in Section VIII-B. Each bar in the figure is divided into encryption latency and I/O latency to illustrate their respective contributions to the total processing time. In the CPU-based implementation, limited concurrency in cryptographic processing and the frequent data transfers between the CPU and host DRAM contribute significantly to the overall latency. In contrast, when cryptographic operations are performed inside in-NAND flash, as implemented in FlashVault, this latency can be mitigated by eliminating the host-side cryptographic processing. Consequently, FlashVault achieves an average latency reduction of 64.2% for decryption and 37.0% for encryption compared to CPU-based processing.
@@ III-B3 Latency of PKC and PQC Verification
Fig. 6 shows the signature verification latency using well-known PKC and PQC during secure boot for boot images from 256 KB to 10 MB. The verification sequence starts by loading the firmware image from boot ROM or NAND flash, computes a cryptographic hash, e.g., SHA [76], and verifies the resulting digest using a digital signature algorithm, e.g., PKC or PQC. The input to the signature verification is relatively small, as it is a fixed-length digest produced by the hash function. However, since boot images range from hundreds of KB to tens of MB, image loading and hash computation account for a large portion of the total verification latency. Meanwhile, the overall verification latency varies with the size of the boot image, increasing as the image size grows. As shown in the figure, when a 10 MB image is used, the signature verification time exceeds 100 ms across all cryptographic algorithms [45]. This result indicates that it is difficult to meet the constraint imposed by commercial hardware implementations, which require boot-time verification to be completed within 100 ms after system power stabilization (i.e., power_good). Moreover, as modern high-end SSDs often adopt firmware images larger than 10 MB [95], meeting the timing constraint using CPU-based verification becomes even more challenging.
@@
AES [96] | 10 / 12 / 14 | 128 / 192 / 256 | 128 | 8 | XOR, S-Box, Shift, Multiplication | ||||||
| 3DES [84] | 48 | 112 / 168 | 64 | 4 | XOR, S-Box, Permutation | |||||||
| IDEA [59] | 8.5 | 128 | 64 | 16 | XOR, ModAdd, ModMult | |||||||
| Serpant [9] | 32 | 128 / 192 / 256 | 128 | 4 | XOR, S-Box, Shift, Permutation | |||||||
| HIGHT [37] | 32 | 128 | 64 | 8 | XOR, Shift, ModAdd, ModMult | |||||||
| SM4 [33] | 32 | 128 | 128 | 8 | XOR, S-Box, Shift | |||||||
| Camellia [4] | 18 / 24 | 128 / 192 / 256 | 128 | 8 | XOR, S-Box, Shift, AND, OR | |||||||
| HASH | SHA-2 [76] | 64 | - | 512 | 32 | XOR, NOT, AND, Shift, ModAdd | ||||||
| PUK | RSA [92] | 1 | 1024 / 2048 / 3072 / 4096 | 256 / 512 | 1024 | AND, OR, Shift, ModMult, ModExp | ||||||
| ECDSA [80] | 1 | 160 / 224 / 256 / 384 / 521 | 256 / 512 | 256 | AND, OR, Shift, ModMult, ModInv | |||||||
| PQC | Dilithium [7, 26] | 1 | 2048 / 3072 | 256 / 512 | 23 | XOR, AND, Shift, NTT, Keccak, ModMult, Comparison | ||||||
| Falcon [89] | 1 | 896 / 1280 | 256 / 512 | 64 | XOR, AND, Shift, FFT, ModMult, Comparison | |||||||
| SPHINCS+ [8] | 1 | 256 / 384 / 512 | 256 / 512 | 32 | XOR, AND, Shift, Keccak, Comparison |
* Round In PKC and PQC, the overall operation is not performed in a repetitive round-based manner, but rather consists of a single large mathematical operation, which is why the round count is set to 1. However, despite having a round count of 1, the internal operations may involve multiple rounds. For instance, in RSA, the modular exponentiation is performed times, depending on the length of the public key exponent [92], whereas in PQC, the Keccak-f1600 permutation function is executed 24 times [63].
@@
Processing Granularity In PUK schemes, the processing granularity represents the bit length of the prime modulus, whereas in PQC, it corresponds to the bit width of polynomial coefficients. Although PQC generally operates on smaller bit-width units than PUK, it achieves comparable security levels through high-degree polynomial operations or hash-based constructions [78].
@@ III-B4 Latency of PKC and PQC Signing
Fig. 7 presents the latency breakdown of digital signature generation using various PKC and PQC algorithms in a tamper-proof logging scenario, where signatures are generated at runtime for individual log entries. Since logging operations typically target small messages ranging from a few hundred bytes to several kilobytes [1], the evaluation was conducted with input message sizes ranging from 256 B to 10 KB. Given such small message sizes, PKC and PQC operations dominate the total latency. Commercial tamper-proof logging systems must complete log generation and signing for each event within 2 to 15 ms to meet real-time processing requirements [99, 1]. However, as shown in the figure, Falcon 1024 and SPHINCS+ require considerable time for signature generation regardless of the input message size. This indicates that several NIST-standardized PQC algorithms are impractical for latency-critical tamper-proof logging when executed on the host CPU, as is typical in modern SSDs.
@@ IV-A Available Silicon Area
A key feature of FlashVault is integrating cryptographic engines into the unused area under the memory array of a 4D V-NAND chip. To enable this integration in the empty space, we first estimate the available area in the 4D V-NAND. Due to lack of public data, we analyze Samsung’s first 4D V-NAND product, the 7th-generation 4D V-NAND chip [14]. We calculate the available space as , where and are the memory array and peripheral circuit areas. To estimate the peripheral circuit area, we refer to the specifications of Samsung’s 6th-generation 3D V-NAND solution [46]. The 7th-generation NAND flash chip improves area efficiency by repositioning the peripheral circuit underneath the memory array. As documented in [31, 18, 38], the peripheral circuit lies entirely under the memory array. Peripheral area is estimated as , where and are die sizes of the 6th and 7th-genration chips. Consequently, we identify a free space of 20.60 in the 7th-generation 4D V-NAND flash. (Takeway 1) We leverage this unused area (20.60 ) to integrate cryptographic engines supporting diverse algorithms.
@@ IV-B Power Budget
Since a NAND flash chip receives only a limited amount of power from the main power management IC of the SSD [126], any cryptographic engine integrated into the unused space of 4D V-NAND must stay within this power budget. If the proposed architecture exceeds this budget, it may degrade cell reliability or performance due to thermal issues [64, 125]. Thus, we identify the power budget and ensure that cryptographic engines operate safely within this limit. However, as Samsung’s 7th-gen V-NAND power budget is undisclosed, we estimate it based on the power budget of the 6th-generation and the reported efficiency improvements in the 7th-generation. Since block-level programming typically consumes the highest power among NAND flash operations [18, 72], we use it as the basis for power budget estimation. For 6th-generation V-NAND, the programming power can be calculated as follows:
where denotes the programming power, is the programming current, and is the maximum operating voltage. Based on technical documents [46, 93], the programming current and maximum operating voltage of the 6th-generation V-NAND are 13.8 mA and 3.6 V, respectively, resulting in a power budget of 49.68 mW. The 7th-generation V-NAND achieves 16% improvement in power efficiency over the 6th generation [104]; thus, its power budget is estimated to be 41.73 mW. (Takeway 2) Given this estimated budget, we design the cryptographic engines to run in parallel under the power constraint (41.73 mW) within the 4D V-NAND die.
@@ IV-C Reconfigurable Cryptographic Engine
One key design goal of FlashVault is supporting diverse cryptographic algorithms through a versatile cryptographic engine. To identify the fundamental cryptographic primitives used across different schemes, we analyzed 17 block ciphers from the Botan library [11]; public-key cryptosystems based on integer factorization and elliptic curve problems; and PQC algorithms, including lattice-based and hash-based schemes, from the liboqs library [81]. Table I summarizes the key primitive operations used in widely adopted block ciphers and public-key algorithms, and NIST-standardized PQC schemes [75]. Our analysis reveals that block ciphers are typically composed of five basic operations: modular arithmetic, logical operations, permutation, substitution (S-Box), and bitwise shift. PKC and PQC algorithms, on the other hand, primarily rely on logical operations, bitwise shifts, number-theoretic transforms, Keccak, modular arithmetic, and comparison operations. (Takeway 3) Based on these observations, we develop a reconfigurable cryptographic engine that supports diverse primitives, enabling FlashVault to efficiently implement various algorithms within the NAND die.
@@ IV-D Error Correction
NAND technology scaling has evolved beyond physical scaling, transitioning from single-level cell (SLC) to triple-level cell (TLC) designs. While multi-bit memory cells significantly increase storage density and reduce cost per bit, this transformation has deteriorated the reliability characteristics of NAND flash, particularly in terms of data retention and endurance. Fig. 8-(a) shows the threshold voltage () distributions of SLC and how they are affected by various error sources. According to previous studies [88, 50, 64, 65, 100], reading or programming a flash cell slightly increases the of adjacent cells by unintentionally injecting electrons into their charge traps (i.e., cell-to-cell interference). In addition, the leakage of electron charges over time decreases the level (i.e., retention loss). As a result of these mechanisms, the level can drift across the reference voltage (), leading to bit errors. As shown in Fig. 8-(b), TLC has narrower voltage windows for each state than SLC, and the technology scaling further shrinks the size of these windows. To ensure data reliability under such conditions, modern SSDs incorporate error correction codes (ECC) within their controllers. Since FlashVault performs cryptographic operations directly within the NAND chip, it is essential to correct bit errors in advance to ensure only error-free data is encrypted. (Takeway 4) To support reliable in-NAND self-encryption, FlashVault integrates error correction capabilities into the NAND flash chip, allowing correction to take place on-chip.
@@ V FlashVault Architecture
4) Cryptographic Unit: The core component of FlashVault is the cryptographic unit, which consists of multiple cryptographic engines. Each engine is configured as an array of block cipher engines (BCEs) for symmetric encryption and an asymmetric cipher engine (ACE) supporting PKC and PQC 333Block ciphers use the same secret key for both encryption and decryption, whereas PKC and PQC use asymmetric key pairs comprising a public key for encryption and a private key for decryption.. When FlashVault operates in self-encrypting mode, block ciphers are executed for every I/O request. These algorithms allow subword-level parallelism across multiple processing stages (i.e., each round in the cipher) [36]. Thus, FlashVault adopts a SIMD-style array of 16 BCEs, enabling high throughput. The ACE module is activated during signature generation and verification. Since most cryptographic operations in PKC and PQC, excluding hash computations, are performed sequentially [55], FlashVault incorporates a single ACE module.
@@
In addition to the cipher engines, the cryptographic engine integrates auxiliary components to support full in-NAND cryptographic functionality. Among them, a low-density parity-check (LDPC) decoder performs error correction on data read from the NAND array. Unlike conventional SSDs, which perform error correction in the SSD controller, FlashVault conducts in-situ correction within the NAND flash chip. For this purpose, the LDPC decoder first corrects raw data errors from the NAND array and then forwards the decoded data to two dual-purpose register files: the Data Register File (DRF) and the Cache Register File (CRF). FlashVault adopts the LDPC decoder in [21], which uses a gradient descent bit-flipping algorithm. It initializes all bits (i.e., variable nodes) using a hard decision, which interprets the signal as 0 or 1. The decoder then iteratively flips the bit with the lowest reliability score, determined from the parity check results and NAND read channel conditions. This process continues until all parity constraints are satisfied or a predefined iteration threshold is reached. Despite being based on a hard-decision architecture with a rate-0.88 quasi-cyclic LDPC code, the decoder provides error correction performance comparable to soft-decision decoders that use probabilistic signal interpretation.
@@
As another auxiliary component, FlashVault includes a Physical Unclonable Function (PUF)-based key management engine (KME). This engine integrates a ring oscillator (RO) PUF that leverages process-induced variations in oscillator frequencies to generate a chip-unique bit pattern, which is used as the root key [67]. The root key is then expanded by a key derivation function (KDF) into symmetric session keys and asymmetric key pairs. The KDF uses a secure hash function, combining the root key with additional salt (i.e., a random value) or context, producing keys with varying lengths and formats, tailored to meet security requirements. The derived keys are stored in the CRF and delivered to either the BCEs or the ACE via the internal router.

### v2

@@ III-B2 Performance of Block Cipher Algorithms
Fig. 5 shows the latency breakdown for six representative block ciphers (3DES [3des], HIGHT [HIGHT], IDEA [idea], Serpent [serpant], SM4 [sm4_report], Camellia [camellia]) executed in CTR mode, measured under both CPU and FlashVault. Although FlashVault embeds multiple cryptographic engines within each NAND flash chip, Fig. 5 reports the result based on a single-block-per-chip configuration to focus on the impact of where encryption occurs. The experimental setup for FlashVault is described in Section VIII-B. Each bar in the figure is divided into encryption latency and I/O latency to illustrate their respective contributions to the total processing time. In the CPU-based implementation, limited concurrency in cryptographic processing and the frequent data transfers between the CPU and host DRAM contribute significantly to the overall latency. In contrast, when cryptographic operations are performed inside in-NAND flash, as implemented in FlashVault, this latency can be mitigated by eliminating the host-side cryptographic processing. Consequently, FlashVault achieves an average latency reduction of 64.2% for decryption and 37.0% for encryption compared to CPU-based processing.
@@ III-B3 Latency of PKC and PQC Verification
Fig. 6 shows the signature verification latency using well-known PKC and PQC during secure boot for boot images from 256 KB to 10 MB. The verification sequence starts by loading the firmware image from boot ROM or NAND flash, computes a cryptographic hash, e.g., SHA [sha], and verifies the resulting digest using a digital signature algorithm, e.g., PKC or PQC. The input to the signature verification is relatively small, as it is a fixed-length digest produced by the hash function. However, since boot images range from hundreds of KB to tens of MB, image loading and hash computation account for a large portion of the total verification latency. Meanwhile, the overall verification latency varies with the size of the boot image, increasing as the image size grows. As shown in the figure, when a 10 MB image is used, the signature verification time exceeds 100 ms across all cryptographic algorithms [secure_boot_100ms]. This result indicates that it is difficult to meet the constraint imposed by commercial hardware implementations, which require boot-time verification to be completed within 100 ms after system power stabilization (i.e., power_good). Moreover, as modern high-end SSDs often adopt firmware images larger than 10 MB [samsung_firmware_size], meeting the timing constraint using CPU-based verification becomes even more challenging.
@@
AES [aes_first] | 10 / 12 / 14 | 128 / 192 / 256 | 128 | 8 | XOR, S-Box, Shift, Multiplication | ||||||
| 3DES [3des] | 48 | 112 / 168 | 64 | 4 | XOR, S-Box, Permutation | |||||||
| IDEA [idea] | 8.5 | 128 | 64 | 16 | XOR, ModAdd, ModMult | |||||||
| Serpant [serpant] | 32 | 128 / 192 / 256 | 128 | 4 | XOR, S-Box, Shift, Permutation | |||||||
| HIGHT [HIGHT] | 32 | 128 | 64 | 8 | XOR, Shift, ModAdd, ModMult | |||||||
| SM4 [sm4_report] | 32 | 128 | 128 | 8 | XOR, S-Box, Shift | |||||||
| Camellia [camellia] | 18 / 24 | 128 / 192 / 256 | 128 | 8 | XOR, S-Box, Shift, AND, OR | |||||||
| HASH | SHA-2 [sha] | 64 | - | 512 | 32 | XOR, NOT, AND, Shift, ModAdd | ||||||
| PUK | RSA [rsa_first] | 1 | 1024 / 2048 / 3072 / 4096 | 256 / 512 | 1024 | AND, OR, Shift, ModMult, ModExp | ||||||
| ECDSA [ecdsa] | 1 | 160 / 224 / 256 / 384 / 521 | 256 / 512 | 256 | AND, OR, Shift, ModMult, ModInv | |||||||
| PQC | Dilithium [dilithium, dilithium_paper] | 1 | 2048 / 3072 | 256 / 512 | 23 | XOR, AND, Shift, NTT, Keccak, ModMult, Comparison | ||||||
| Falcon [falcon_first] | 1 | 896 / 1280 | 256 / 512 | 64 | XOR, AND, Shift, FFT, ModMult, Comparison | |||||||
| SPHINCS+ [sphincs] | 1 | 256 / 384 / 512 | 256 / 512 | 32 | XOR, AND, Shift, Keccak, Comparison |
* Round In PKC and PQC, the overall operation is not performed in a repetitive round-based manner, but rather consists of a single large mathematical operation, which is why the round count is set to 1. However, despite having a round count of 1, the internal operations may involve multiple rounds. For instance, in RSA, the modular exponentiation is performed times, depending on the length of the public key exponent [rsa_first], whereas in PQC, the Keccak-f1600 permutation function is executed 24 times [keccak_permutation].
@@
Processing Granularity In PUK schemes, the processing granularity represents the bit length of the prime modulus, whereas in PQC, it corresponds to the bit width of polynomial coefficients. Although PQC generally operates on smaller bit-width units than PUK, it achieves comparable security levels through high-degree polynomial operations or hash-based constructions [pqc_small_length].
@@ III-B4 Latency of PKC and PQC Signing
Fig. 7 presents the latency breakdown of digital signature generation using various PKC and PQC algorithms in a tamper-proof logging scenario, where signatures are generated at runtime for individual log entries. Since logging operations typically target small messages ranging from a few hundred bytes to several kilobytes [tamper_logging_size], the evaluation was conducted with input message sizes ranging from 256 B to 10 KB. Given such small message sizes, PKC and PQC operations dominate the total latency. Commercial tamper-proof logging systems must complete log generation and signing for each event within 2 to 15 ms to meet real-time processing requirements [emlog_logging, tamper_logging_size]. However, as shown in the figure, Falcon 1024 and SPHINCS+ require considerable time for signature generation regardless of the input message size. This indicates that several NIST-standardized PQC algorithms are impractical for latency-critical tamper-proof logging when executed on the host CPU, as is typical in modern SSDs.
@@ IV-A Available Silicon Area
A key feature of FlashVault is integrating cryptographic engines into the unused area under the memory array of a 4D V-NAND chip. To enable this integration in the empty space, we first estimate the available area in the 4D V-NAND. Due to lack of public data, we analyze Samsung’s first 4D V-NAND product, the 7th-generation 4D V-NAND chip [samsung_7]. We calculate the available space as , where and are the memory array and peripheral circuit areas. To estimate the peripheral circuit area, we refer to the specifications of Samsung’s 6th-generation 3D V-NAND solution [samsung_6]. The 7th-generation NAND flash chip improves area efficiency by repositioning the peripheral circuit underneath the memory array. As documented in [4d_nand_survey, pif, CUA_patent], the peripheral circuit lies entirely under the memory array. Peripheral area is estimated as , where and are die sizes of the 6th and 7th-genration chips. Consequently, we identify a free space of 20.60 in the 7th-generation 4D V-NAND flash. (Takeway 1) We leverage this unused area (20.60 ) to integrate cryptographic engines supporting diverse algorithms.
@@ IV-B Power Budget
Since a NAND flash chip receives only a limited amount of power from the main power management IC of the SSD [power_supply], any cryptographic engine integrated into the unused space of 4D V-NAND must stay within this power budget. If the proposed architecture exceeds this budget, it may degrade cell reliability or performance due to thermal issues [heatwatch, thermal_performance]. Thus, we identify the power budget and ensure that cryptographic engines operate safely within this limit. However, as Samsung’s 7th-gen V-NAND power budget is undisclosed, we estimate it based on the power budget of the 6th-generation and the reported efficiency improvements in the 7th-generation. Since block-level programming typically consumes the highest power among NAND flash operations [pif, power_dominant_program], we use it as the basis for power budget estimation. For 6th-generation V-NAND, the programming power can be calculated as follows:
where denotes the programming power, is the programming current, and is the maximum operating voltage. Based on technical documents [samsung_6, samng_current], the programming current and maximum operating voltage of the 6th-generation V-NAND are 13.8 mA and 3.6 V, respectively, resulting in a power budget of 49.68 mW. The 7th-generation V-NAND achieves 16% improvement in power efficiency over the 6th generation [power_efficiency_improve]; thus, its power budget is estimated to be 41.73 mW. (Takeway 2) Given this estimated budget, we design the cryptographic engines to run in parallel under the power constraint (41.73 mW) within the 4D V-NAND die.
@@ IV-C Reconfigurable Cryptographic Engine
One key design goal of FlashVault is supporting diverse cryptographic algorithms through a versatile cryptographic engine. To identify the fundamental cryptographic primitives used across different schemes, we analyzed 17 block ciphers from the Botan library [botan]; public-key cryptosystems based on integer factorization and elliptic curve problems; and PQC algorithms, including lattice-based and hash-based schemes, from the liboqs library [liboqs]. Table I summarizes the key primitive operations used in widely adopted block ciphers and public-key algorithms, and NIST-standardized PQC schemes [nist_pqc_altorithms]. Our analysis reveals that block ciphers are typically composed of five basic operations: modular arithmetic, logical operations, permutation, substitution (S-Box), and bitwise shift. PKC and PQC algorithms, on the other hand, primarily rely on logical operations, bitwise shifts, number-theoretic transforms, Keccak, modular arithmetic, and comparison operations. (Takeway 3) Based on these observations, we develop a reconfigurable cryptographic engine that supports diverse primitives, enabling FlashVault to efficiently implement various algorithms within the NAND die.
@@ IV-D Error Correction
NAND technology scaling has evolved beyond physical scaling, transitioning from single-level cell (SLC) to triple-level cell (TLC) designs. While multi-bit memory cells significantly increase storage density and reduce cost per bit, this transformation has deteriorated the reliability characteristics of NAND flash, particularly in terms of data retention and endurance. Fig. 8-(a) shows the threshold voltage () distributions of SLC and how they are affected by various error sources. According to previous studies [modern_ssd_asplos, evanesco, heatwatch, nand_pomacs, 3d_nand_reliability_nand], reading or programming a flash cell slightly increases the of adjacent cells by unintentionally injecting electrons into their charge traps (i.e., cell-to-cell interference). In addition, the leakage of electron charges over time decreases the level (i.e., retention loss). As a result of these mechanisms, the level can drift across the reference voltage (), leading to bit errors. As shown in Fig. 8-(b), TLC has narrower voltage windows for each state than SLC, and the technology scaling further shrinks the size of these windows. To ensure data reliability under such conditions, modern SSDs incorporate error correction codes (ECC) within their controllers. Since FlashVault performs cryptographic operations directly within the NAND chip, it is essential to correct bit errors in advance to ensure only error-free data is encrypted. (Takeway 4) To support reliable in-NAND self-encryption, FlashVault integrates error correction capabilities into the NAND flash chip, allowing correction to take place on-chip.
@@ V FlashVault Architecture
4) Cryptographic Unit: The core component of FlashVault is the cryptographic unit, which consists of multiple cryptographic engines. Each engine is configured as an array of block cipher engines (BCEs) for symmetric encryption and an asymmetric cipher engine (ACE) supporting PKC and PQC 333Block ciphers use the same secret key for both encryption and decryption, whereas PKC and PQC use asymmetric key pairs comprising a public key for encryption and a private key for decryption.. When FlashVault operates in self-encrypting mode, block ciphers are executed for every I/O request. These algorithms allow subword-level parallelism across multiple processing stages (i.e., each round in the cipher) [block_cipher_review]. Thus, FlashVault adopts a SIMD-style array of 16 BCEs, enabling high throughput. The ACE module is activated during signature generation and verification. Since most cryptographic operations in PKC and PQC, excluding hash computations, are performed sequentially [survey_pkc], FlashVault incorporates a single ACE module.
@@
In addition to the cipher engines, the cryptographic engine integrates auxiliary components to support full in-NAND cryptographic functionality. Among them, a low-density parity-check (LDPC) decoder performs error correction on data read from the NAND array. Unlike conventional SSDs, which perform error correction in the SSD controller, FlashVault conducts in-situ correction within the NAND flash chip. For this purpose, the LDPC decoder first corrects raw data errors from the NAND array and then forwards the decoded data to two dual-purpose register files: the Data Register File (DRF) and the Cache Register File (CRF). FlashVault adopts the LDPC decoder in [ldpc], which uses a gradient descent bit-flipping algorithm. It initializes all bits (i.e., variable nodes) using a hard decision, which interprets the signal as 0 or 1. The decoder then iteratively flips the bit with the lowest reliability score, determined from the parity check results and NAND read channel conditions. This process continues until all parity constraints are satisfied or a predefined iteration threshold is reached. Despite being based on a hard-decision architecture with a rate-0.88 quasi-cyclic LDPC code, the decoder provides error correction performance comparable to soft-decision decoders that use probabilistic signal interpretation.
@@
As another auxiliary component, FlashVault includes a Physical Unclonable Function (PUF)-based key management engine (KME). This engine integrates a ring oscillator (RO) PUF that leverages process-induced variations in oscillator frequencies to generate a chip-unique bit pattern, which is used as the root key [ro_puf]. The root key is then expanded by a key derivation function (KDF) into symmetric session keys and asymmetric key pairs. The KDF uses a secure hash function, combining the root key with additional salt (i.e., a random value) or context, producing keys with varying lengths and formats, tailored to meet security requirements. The derived keys are stored in the CRF and delivered to either the BCEs or the ACE via the internal router.